the general form of the constant function is : f(x)=a(x^0), where a is constant.
So when the function passes by any point on the y-axis, the function will be :f(x)=a(0^0).
This means there must be a hole in the point of intersection of the y-axis and the constant function on the graph because (0^0) is undefined.
So why don't we draw the hole on the Cartesian coordinate plane ?

Answer by stanbon(75887)

the general form of the constant function is : f(x)=a(x^0), where a is constant.
So when the function passes by any point on the y-axis, the function will be :f(x)=a(0^0).
This means there must be a hole in the point of intersection of the y-axis and the constant function on the graph because (0^0) is undefined.
So why don't we draw the hole on the Cartesian coordinate plane ?
-----
Any number to the zero power is 1.
So f(x) = a(x^0) = a
The y-intercept is (0,a)
==========================
Cheers,
Stan H.
=======================